Ken needs to estimate the beta for his firm’s pharmacy division to ultimately calculate a WACC for this division. The division has a: weight of equity = 80.0%; weight of debt = 20.0%; and tax rate = 13.5%. Ken has identified 3 peer firms and found their betas, tax rates, and weights of debt and equity:

Company Beta Tax Rate Weight of Equity Weight of Debt
A 0.99 19.8% 54.8% 45.2%
B 1.75 13% 12.6% 87.4%
C 1.66 16.3% 64.7% 35.3%

Ken knows that to estimate a divisional beta he will have to:

1. Unlever the betas of the 3 peer firms using equation 15-12 (in the text and the module 8 supplement)

2. Take an average of the 3 unlevered betas.

3. Re-lever the average beta based on the division’s weights of equity and debt as well as the tax rate.

Based on the information above, what is the beta that Ken should use for the beta for his division?
Answer #1

Unlevered Beta = Levered Beta / [1 + {(1 - t) * (D/E)}]

A's Unlevered Beta = 0.99 / [1 + {(1 - 0.198) * (45.2/54.8)}] = 0.99 / [1 + 0.6615] = 0.5958

B's Unlevered Beta = 1.75 / [1 + {(1 - 0.13) * (87.4/12.6)}] = 1.75 / [1 + 6.0348] = 0.2488

C's Unlevered Beta = 1.66 / [1 + {(1 - 0.163) * (35.3/64.7)}] = 1.66 / [1 + 0.4567] = 1.1396

Average Unlevered Beta = [0.5958 + 0.2488 + 1.1396] / 3 = 1.9842 / 3 = 0.6614

Ken's Levered Beta = Unlevered Beta * [1 + {(1 - t) * (D/E)}]

= 0.6614 * [1 + {(1 - 0.135) * (20/80)}] = 0.6614 * [1 + 0.21625] = 0.8044, or 0.80